How do I factor thr trinomial for
16u^2+8u-35

Answer by scott8148(6628)

16 = 4 * 4 ___ -35 = -5 * 7

4 * -5 = -20 ___ 4 * 7 = 28

-20 + 28 = 8

(4u + 7)(4u - 5)